OpenStudy (anonymous):

3x-2 greater than or equal to -7 explain.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@DanJS

OpenStudy (dtan5457):

\[3x-2\ge-7 \]

OpenStudy (dtan5457):

add the two to both sides

OpenStudy (dtan5457):

divide by 3

OpenStudy (dtan5457):

\[x \ge \frac{ -5 }{ 3 }\]

OpenStudy (anonymous):

okay so for -9x-3 greater than or equal to 51 i add 3 to both sides?

OpenStudy (anonymous):

@bibby ^

Nnesha (nnesha):

yes right

Nnesha (nnesha):

first add 3 both side then divide both side by -9 and remember when you divide or multiply by negative number you must flip the sign

OpenStudy (anonymous):

-9x-6 greater than or equal to 54??

Nnesha (nnesha):

when you add 3 on left side so then you should have -3 +3 so its equal to what ???

OpenStudy (anonymous):

i really dont know

Nnesha (nnesha):

okay so what is -3 +3??

Nnesha (nnesha):

 now add 3 both side